Circular No. 6906

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1998 J1 (SOHO) More than 100 positions were reported (cf. IAUC 6894, MPEC 1998-

J13, -J14 and -J21) of this comet as it moved across the top of the field of the SOHO-LASCO C3 coronagraph between May 3.78 and 8.87 UT. Attempts to record the comet from the ground were often inconclusive, mainly because of the difficulty of predicting the comet’s position from the rough SOHO data. Visual observations that appear to be positive are: May 11.23 UT, m_{1} ∼ 0.5, coma diameter D = 0′.5 (N. Biver, Kahe Point, Oahu, HI, 0.26-m reflector); 11.23, m_{1} ∼ 0.5, D = 1′.0 (H. Dahle, same location, 20×80 binoculars); 14.35, m_{1} ∼ 2–3, tail ∼ 15′ long toward the east (P. Nation, Christies Beach, South Australia, 11×80 binoculars); 16.24, m_{1} ∼ 3.5, D = 1′ and rather condensed with faint, short, fanshaped tail ∼ 5′ long in p.a. 110^{o} (Biver); 16.34, m_{1} ∼ 2.5–3, tail > 0^{o}.5 long in p.a. 90–130^{o} (C. E. Drescher, Warrill View, Queensland, Australia, 7 × 50 binoculars). In addition, the comet was apparently imaged (focal length 500 mm, ST-6) by O. Farago (Stuttgart, Germany) on May 10.80 UT, when the close proximity to the Pleiades allowed J. Jahn (Bodenteich, Germany) to perform astrometry (internal consistency ∼ 2′′) from four stars bright enough for detection in strong twilight at 4^{o}.5 altitude in the 1^{o} field! Positions and a compromise parabolic orbit and ephemeris follow:

1998 UT α_{2000} δ_{2000} Observer
May 10.80208 3^{h}45^{m}03.^{s}6 +23^{o}43′15′′ Farago
16.249  4 47.7  +12 53  Biver
16.34  4 48  +12 50  Drescher
T = 1998 May 8.352 TT  ω = 107.^{o}649
Ω = 355.021 }2000.0
q = 0.15015 AU  i = 68.346
1998TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
May 13 4^{h}13.^{m}70 +19^{o}54.^{′}2 0.894 0.246 13.^{o}1 111.^{o}6 1.7
15 4 35.43 +15 41.6 0.865 0.309 16.7 109.5 2.6
17 4 54.17 +11 18.8 0.851 0.371 20.8 104.8 3.3
19 5 10.73 + 7 00.0 0.848 0.432 24.9  99.3 4.0
21 5 25.62 + 2 52.4 0.853 0.490 28.9  93.8 4.6
23 5 39.18 − 1 00.0 0.865 0.547 32.7  88.6 5.1
25 5 51.67 − 4 35.6 0.883 0.603 36.3  83.7 5.5
27 6 03.27 − 7 54.1 0.905 0.656 39.5  79.2 6.0
1998 May 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden